Consider the reaction below

I2O5(g) + 5 CO2(g) ———> 5 CO2(g) + I2(g)


80.0g of iodine (v) oxide reacts with 28.0g of carbon monoxide. Determine the mass of iodine I2 which would be produced


1
Expert's answer
2022-03-17T14:58:53-0400

Solution:

The molar mass of iodine(V) oxide (I2O5) is 333.81 g/mol

The molar mass of carbon monoxide (CO) is 28.01 g/mol


Calculate moles of each reactant:

Moles of I2O5 = (80.0 g I2O5) × (1 mol I2O5 / 333.81 g I2O5) = 0.23966 mol I2O5

Moles of CO = (28.0 g CO) × (1 mol CO / 28.01 g CO) = 0.99964 mol CO


Balanced chemical equation:

I2O5(g) + 5CO(g) ⟶ 5CO2(g) + I2(g)

According to stoichiometry:

1 mol of I2O5 reacts with 5 mol of CO

Thus, 0.23966 mol I2O5 reacts with:

(0.23966 mol I2O5) × (5 mol CO / 1 mol I2O5) = 1.1983 mol CO

However, initially there is 0.99964 mol of CO (according to the task).

Therefore, CO acts as limiting reactant and I2O5 is excess reactant.


According to stoichiometry:

5 mol of CO produces 1 mol of I2

Thus, 0.99964 mol of CO produces:

( 0.99964 mol CO) × (1 mol I2 / 5 mol CO) = 0.19993 mol I2


The molar mass of iodine (I2) is 253.809 g/mol

Therefore,

Mass of I2 = (0.19993 mol I2) × (253.809 g I2 / 1 mol I2) = 50.744 g I2 = 50.7 g I2

Mass of I2 = 50.7 g


Answer: 50.7 grams of iodine (I2) would be produced